OpenStudy (anonymous):

What is the sum of the two vectors (2,-1) and (3,3)?

OpenStudy (anonymous):

vector addition is just (a1+b1,a2+b2), so in your case, it's (5,2)

OpenStudy (stamp):

Va = 2i - j Vb = 3i + 3j Vr = Va + Vb Find Vr.

OpenStudy (anonymous):

\[\left(\begin{matrix}2 \\ -1\end{matrix}\right) +\left(\begin{matrix}3 \\ 3\end{matrix}\right)=\left(\begin{matrix}5 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

Thank You @thushananth01 @stamp &'d @johnny0929